20. The density of liquid water is 0.997 g/mL at 25°C.

a. Calculate the mass of 250.0 mL (about a cupful) of water.

b. How many moles of water are in 250.0 mL of water? Hint: Use the result of

(a).

c. Calculate the volume that would be occupied by 2.000 mol of water at 25°C.

d. What mass of water is 2.000 mol of water?


1
Expert's answer
2020-10-06T09:16:57-0400

a. Calculate the mass of 250.0 mL:


"m=\rho V=0.997\cdot250=249.3\text{ g}."


b. 250.0 mL of water contain the following amount of moles:


"n=\frac{m}{\Mu}=\frac{249.25\text{ g}}{18\text{ g/mol}}=13.85\text{ mol}."


c. The volume that would be occupied by 2.000 (two) mol of water at 25°C is


"V=\frac{m}{\rho}=\frac{n\Mu}{\rho}=\frac{2\cdot18}{0.997}=36.11\text{ mL}."


d. Finally, find what mass of water is 2.000 (two) mol of water:


"m=n\Mu=2\text{ mol}\cdot18\text{ g/mol}=36\text{ g}."